# Sensing with near-infrared laser trapped fluorescent nanodiamonds

### v2

@@ V Discussion
The shift in the ODMR resonance frequency with increasing NIR power, as shown in Fig. 4, Fig. 6, and Fig. 7a), suggests an increase in temperature within the FNDs. This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K [Acosta2010, Toyli2012]. However, the extent of the thermal effect is significantly influenced by the surrounding medium’s properties. Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs. This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination. Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power, as shown in Fig. 6 and Fig. 7.
In addition to thermal effects, the NIR laser alters the charge-state dynamics of the NV centers. Our simulations of charge-state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons. However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases. This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
@@
The total fluorescence intensity decreases with increasing NIR power, as shown in Fig. 9. This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [Chapman2010]. The increase in $\sigma_{0-}$ with higher NIR power, depicted in Fig. 6 and Fig. 7, further confirms the shift in charge state distribution. Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations. This is reflected in the decreased ODMR contrast observed in our measurements (Fig. 4). Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times [Aslam2013, Goudarzi2019].
@@
While thermal effects and charge-state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered. Multi-photon absorption processes could directly excite the NV centers without significant heating, yet, the relationship between ODMR frequency shift and NIR power (Fig. 6) is contrary to this explanation. Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium. The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, while those in FNDs trapped in water are not.
In our opinion, the data presented in this work establish a clear hierarchy of influences: (i) the thermal conductivity and heat capacity of the surrounding medium dictate the magnitude of the NIR-induced temperature rise and therefore the thermally induced shifts in ODMR frequency and the shortening of $T_{\rm relax}$; (ii) the NIR photon flux sets the rate of photo-ionisation and recombination, which in turn controls the NV0/NV- balance and the fluorescence contrast. Optical trapping itself plays a mainly mechanical role-holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
@@
The results presented in Section IV confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds. In particular, Fig. 10a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
VI Conclusion
This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities. By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and fluorescence relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
@@ VI Conclusion
However, this study also reveals critical uncertainties and challenges that must be addressed to fully exploit the potential of NV-based FNDs. Long-term stability under extended NIR irradiation remains an open question, particularly regarding potential degradation of fluorescence and spin properties. The effects of complex biological environments, including interactions with proteins, ions, and other cellular components, are yet to be fully understood. The role of surface functionalization in enhancing NV center stability and sensing accuracy requires further exploration, as does the potential for alternative NIR wavelengths or advanced laser modulation techniques to improve performance.
Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications. Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a a real-time sensing protocol inside cells by using optically trap FNDs in the cellular medium. Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step in translating this research into practical diagnostic tools.
@@ acknowledgments
We specify in Table 2 all part numbers in the optical setup, Fig. 2a).
Table 2: Optical setup component
@@ 7 Iris Aperture ThorLabs ID12Z/M
8 IR disk ThorLabs VRC4CPT
9 Beam expander
10 Right-angle prism ThorLabs DFM1/M
@@ 20 Lens ThorLabs LA1509 - A
21 Right-angle prism ThorLabs DFM1/M
@@ 22 Lens ThorLabs LA1484 - A
23 Filters Bightline Basic 460/60 - ThorLabs MF630-69 Emission filter
@@ 25 Filter ThorLabs NE20A
26 Filters Bightline Basic FF01-630/52 - ThorLabs FELH 0550
@@ 34 Quadrant photoelectric detector ThorLabs PDQ80A
35 White lamp
@@ 36 Spectrometer Ocean Optics QE Pro
In literature, a varying range of values for the rate constants used in the numerical model corresponding to Fig. 8 exist. We applied the relevant values from Ref. [Qian2022] and combined them with parameters from [Meirzada2018] for the four ionization and recombination rates as these authors use the same laser wavelengths as us in their protocol. Below, in Table 3, we provide the values used in the simulations of Section III.4.
@@
laser_green_power 0.7 mW Green laser power

### v3

@@ 4 Discussion
The shift in the ODMR resonance frequency with increasing NIR power, as shown in Fig. 3a), and Fig. 4, suggests an increase in temperature within the FNDs. This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K Acosta et al. (2010); Toyli et al. (2012) and corresponds with previous suggestions of the heating effect of NIR light Lai et al. (2013). However, the extent of the thermal effect is significantly influenced by the surrounding medium’s properties. Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs. This is supported by our temperature simulations, which show much lower temperature rise in water compared to air during NIR illumination. Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power.
@@
In addition to thermal effects, the NIR laser alters the charge-state dynamics of the NV centers. Our simulations of charge-state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons. However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases. This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers. As a result, the total fluorescence intensity decreases with increasing NIR power, as shown in Fig. 5. This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers Chapman and Plakhotnik (2011). The increase in $\sigma_{0-}$ with higher NIR power, depicted in Fig. 3b) and Fig. 4, further confirms the shift in charge state distribution. Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations. This is reflected in the decreased ODMR contrast observed in our measurements (data not shown). Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times Aslam et al. (2013). With an all-optical protocol for relaxometry as used here, the results for $T_{\rm relax}$ presented inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states. To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed. Examples include the microwave $\pi$-pulse method introduced by Ref. Jarmola et al. (2012), as well as alternative sequences described in Ref. Cardoso Barbosa et al. (2023).
@@
While thermal effects and charge-state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered. Multi-photon absorption processes could directly excite the NV centers without significant heating, yet, the relationship between ODMR frequency shift and NIR power (Fig. 4) is contrary to this explanation. Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium. The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, while those in FNDs trapped in water are not. It should also be emphasized that while our results are consistent with an interpretation of ODMR frequency shift being proportional to a temperature shift, the interpretation may be more complex since changes in local electrical field also implies shifts of the ODMR frequency Dolde et al. (2011), an effect that we cannot completely exclude in our experiments. Similarly, literature suggests NIR-induced dark state transitions with strong NIR pulses Geiselmann et al. (2013); Cao et al. (2025b, a), and including these extra states in our theoretical model could provide more detailed understanding of our results.
In our opinion, the data presented in this work establish a clear hierarchy of influences: (i) the thermal conductivity and heat capacity of the surrounding medium dictate the magnitude of the NIR-induced temperature rise and therefore the thermally induced shifts in ODMR frequency and the shortening of $T_{\rm relax}$; (ii) the NIR photon flux sets the rate of photo-ionisation and recombination, which in turn controls the NV0/NV- balance and the fluorescence contrast. Optical trapping itself plays a mainly mechanical role–holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
@@
The results presented in the sensing section confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds. In particular, Fig. 6a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
5 Conclusion
This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive overview of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities. By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and fluorescence relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
@@ 5 Conclusion
However, this study also reveals critical uncertainties and challenges that must be addressed to fully exploit the potential of NV-based FNDs. Long-term stability under extended NIR irradiation remains an open question, particularly regarding potential degradation of fluorescence and spin properties. The role of surface functionalization in enhancing NV center stability and sensing accuracy requires further exploration, as does the potential for alternative NIR wavelengths or advanced laser modulation techniques to improve performance. The effects of complex biological environments, including interactions with proteins, ions, and other cellular components, are yet to be fully understood.
Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications. Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a a real-time sensing protocol inside cells by using optically trap FNDs in the cellular medium.
@@ Acknowledgments
Figure 9: Detailed illustration of the optical setup with the most important lasers, optical lenses, filters and equipment. The components are specified in Table 2.
We specify in Table 2 all part numbers in the optical setup, Fig. 9.
Table 2: Components of the optical setup, Fig. 9. Filters at position 23 and 26 are switchable.
@@ 7 Iris Aperture ThorLabs ID12Z/M
9 Beam expander Part of Thorlabs OTKB/M
10 Dicroic mirror DMSP805R (in ThorLabs DFM1/M)
@@ 20 Lens ThorLabs LA1509 - A
21 Dicroic mirror Semrock NFD01-532-25x36 (in Thorlabs DFM1/M)
@@ 22 Lens ThorLabs LA1484 - A
23 Filters Brightline Basic 460/60 - ThorLabs MF630-69 Emission filter
@@ 25 Filter ThorLabs NE20A
26 Filters Brightline Basic FF01-630/52 - ThorLabs FELH 0650
@@ 34 Quadrant photoelectric detector ThorLabs PDQ80A
35 White lamp Part of Thorlabs OTKB/M
@@ 36 Spectrometer Ocean Optics QE Pro
In literature, a varying range of values for the rate constants used in the numerical model corresponding to Fig. 2 exist. We applied the relevant values from Ref. Qian et al. (2022) and combined them with parameters from Meirzada et al. (2018) for the four ionization and recombination rates as these authors use the same laser wavelengths as us in their protocol. Below, in Table 3, we provide the values used in the simulations of Section 2.5.
@@
$I_{G}$ 0.7 mW Green laser power